OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? a. 2 b.-4/5 c.5/4 d.-2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use slope formula : http://www.algebra-class.com/image-files/slope-formula-2.gif

OpenStudy (anonymous):

idk how

ganeshie8 (ganeshie8):

(1, -5) and (4, 1) x1 y1 x2 y2

ganeshie8 (ganeshie8):

slope = \(\large \dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

a.\[\frac{1--5}{4-1}=2 \]

OpenStudy (anonymous):

the answer is A.